Two cards are drawn from the pack of 52 cards. Find the probability that both are aces


1
Expert's answer
2021-10-19T13:49:48-0400

desired probability "= \frac{4}{52} \times \frac{3}{51}= \frac{12}{2652} = 0.00452"